Blending spices. Spice of Life sells ground sumac for $1.35 an ounce and ground thyme for $1.85 an ounce. Aman wants to make a 20-oz Zahtar seasoning blend using the two spices that sells for $1.65 an ounce. How much of each spice should Aman use?

Answer by htmentor(1343)

Let x = the amount of sumac
Then 20 - x = the amount of thyme
The price of the blend is determined by the price of the seasonings added together:
1.65*20 = 1.35x + 1.85(20-x)
20(1.65-1.85) = -0.5x
This gives x = 8
Therefore, he should use 8 oz. of sumac, and 12 oz. of thyme.